# In celebration of Prof. Xun Yu Zhou’s 60th birthday

### v1

@@ 5 Stability of the Solutions to the SBP
We note that all the results in the previous section are based on an important assumption: ${\mu_{\rm ini}}=\delta_{x_{0}}$, for some $x_{0}\in\mathbb{R}^{d}$. In this and the next section, we shall extend the results to more general initial condition ${\mu_{\rm ini}}\in\mathscr{P}_{2}(\mathbb{R}^{d})$, and establish a similar rate of convergence.
@@
For any $\mu_{0},\mu_{T}\in\mathscr{P}(\mathbb{R}^{d})$, there exists a unique pair of $\sigma$-finite measures $\nu_{0},\nu_{T}\in\mathscr{M}(\mathbb{R}^{d})$ such that the measure $\pi\in\mathscr{P}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ defined by
@@
has marginals $\mu_{0}$ and $\mu_{T}$. Furthermore, $\nu_{T}$ and $\mu_{T}$ (resp. $\nu_{0}$ and $\mu_{0}$) are mutually absolutely continuous, denoted by $\nu_{T}\simeq\mu_{T}$ (resp. $\nu_{0}\simeq\mu_{0}$). ∎
Following Proposition 5.1, let us denote a (well-defined) mapping ${\cal T}:\mathscr{P}_{2}(\mathbb{R}^{d})\times\mathscr{P}_{2}(\mathbb{R}^{d})\mapsto\mathscr{M}(\mathbb{R}^{d})\times\mathscr{M}(\mathbb{R}^{d})$ by ${\cal T}(\mu_{0},\mu_{T})=(\nu_{0},\nu_{T})$. In particular, in what follows we shall often fix $\mu_{0}={\mu_{\rm ini}}\in\mathscr{P}_{2}(\mathbb{R}^{d})$, and focus mainly on $\nu_{T}$. Note that in Proposition 5.1 the measures $(\nu_{0},\nu_{T})$ are only $\sigma$-finite in general, to facilitate our discussion, we shall consider, for a given $\mu_{0}$, the following set:
$\mathscr{D}_{\mu_{0}}:=\{\mu\in\mathscr{P}_{2}(\mathbb{R}^{d}):{\cal T}(\mu_{0},\mu)\ll Leb(\cdot);\,{\cal T}(\mu_{0},\mu)(\mathbb{R}^{d}\times\mathbb{R}^{d})<\infty\}.$
Here $Leb(\cdot)$ denotes the Lebesgue measure on $\mathbb{R}^{d}\times\mathbb{R}^{d}$. In the case when $\mu_{0}={\mu_{\rm ini}}$ is fixed in the discussion, we shall simply denote $\mathscr{D}=\mathscr{D}_{{\mu_{\rm ini}}}$ when context is clear.
We note that if $\mu\in\mathscr{D}$ and $(\nu_{0},\nu_{T})={\cal T}({\mu_{\rm ini}},\mu)$, then $\nu_{T}$ must have a density function, which we shall denote by $\rho^{\mu}\in\mathbb{L}^{1}(\mathbb{R}^{d})$. Moreover, we define an operator $S:\mathscr{P}_{2}(\mathbb{R}^{d})\to\mathscr{P}_{2}(\mathbb{R}^{d})$ by
@@
Moreover, for fixed ${\mu_{\rm ini}}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ and $\mu\in\mathscr{D}=\mathscr{D}_{{\mu_{\rm ini}}}$, we define $h^{\mu}(t,x)=\int_{\mathbb{R}^{d}}p(T,z;t,x)\rho^{\mu}(z)dz$. Then, we have the following result.
@@
Let $\mu\in\mathscr{D}=\mathscr{D}_{{\mu_{\rm ini}}}$. Then, for any $\{\alpha_{t}\}\in\mathbb{L}^{2}_{\mathbb{F}^{0}}([0,T])$, it holds that
@@
From Proposition 5.2 and Lemma 5.3 we see that the density function $\rho^{\mu}$ plays an important role in the structure of the solution of SBP. We shall be particularly interested in the continuous dependence of $\rho^{\mu}:=\Gamma_{1}(\mu)$ on $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$, which we shall refer to as the Stability of the SBP, borrowing the well-known concept of the SBP theory (cf. e.g., [46, 22, 10]). In light of (6.4), we see that if both $\mu_{\rm ini}$} and $\mu$ have densities, then so does ${\cal T}({\mu_{\rm ini}},\mu)$. Furthermore, in light of (5.3), for any $\rho\in\mathbb{L}^{1}(\mathbb{R}^{d})$, we define $h^{\rho}(t,x)=\int_{\mathbb{R}^{d}}p(T,z;t,x)\rho(z)dz$. Then clearly we have $h^{\mu}(t,x)=h^{\rho^{\mu}}(t,x)$.
@@
There exists a function $g\in\mathbb{L}^{2}(\mathbb{R}^{d};(0,1])$, with $\int_{\mathbb{R}^{d}}|x|^{2}g(x)ds<\infty$, and a constant $K>0$, such that $\Big\|\frac{f_{\mu}}{g^{2}}\Big\|_{\infty}\leq K$, for all $\mu\in\mathcal{K}$.
@@
A typical example of the function $g$ in Assumption 5.4 is $e^{-c|x|}$ or $e^{-c|x|^{2}}$, $x\in\mathbb{R}^{d}$, $c>0$. In such a case we see that part (i) holds for all $p>0$. Assumption 5.4 amounts to saying that we focus only on those density functions that have a similar decay rate to function $g$ at $x\sim\infty$. In fact, in light of the estimate (3.2), such a property holds essentially for all transition probabilities of diffusion processes. ∎
@@
$f^{m}_{\mu_{n}}(x)=[\varphi^{m}*f_{\mu_{n}}](y)=\int_{\mathbb{R}^{d}}\varphi^{m}(x-z)f_{\mu_{n}}(z)dz;\quad f^{m}_{\mu}(x)=[\varphi^{m}*f_{\mu}](x),\quad x\in\mathbb{R}^{d}.$
@@
$\leq$ $K\int_{\mathbb{R}^{d}}|\varphi^{m}(z^{\prime}-y)-\varphi^{m}(z^{\prime})|dz^{\prime}\to 0,\quad\mbox{as $y\to 0$.}$
@@
We are now ready to study our main stability result. More precisely, we shall argue that the mapping $\Gamma_{1}:\mathscr{P}_{2}(\mathbb{R}^{d})\mapsto\mathbb{L}^{1}(\mathbb{R}^{d})$ is continuous. That is, that $\mu_{n}$ weakly converges to $\mu_{\rm tar}$} in Prohorov metric would imply that $\rho^{\mu_{n}}$ converges to $\rho^{{\mu_{\rm tar}}}$ in $\mathbb{L}^{1}(\mathbb{R}^{d})$. Such a result, to the best of our knowledge, is novel in the literature.
To simplify our discussion, in what follows, we assume $T=1$, and denote the measure $\mu$ in (5.1) by $\pi$ for notational clarity. Recall that $\pi$ has marginals $\mu_{\rm ini}$} and $\mu$, and in what follows, we shall assume that $\mu_{\rm ini}$} is fixed and $\mu\in\mathscr{E}$. Let us now consider the following entropic optimal transport problem:
@@
$\widehat{\pi}(dxdy)=p(1,y;0,x)e^{\phi^{\mu}(x)+\psi^{\mu}(y)}f_{0}(x)f_{\mu}(y)dxdy,$
where $\phi^{\mu}$ and $\psi^{\mu}$ are the Schrödinger potentials, and we shall enforce the symmetric normalization so that they satisfy (5.11). Since $\widehat{\pi}$ has the marginals $\mu_{\rm ini}$} and $\mu$, by the uniqueness of $(\nu_{0},\nu)={\cal T}({\mu_{\rm ini}},\mu)$, whence $(\rho_{0},\rho^{\mu})$, in Proposition 5.1, we can conclude that
$p(1,y;0,x)\rho_{0}(x)\rho^{\mu}(y)=p(1,y;0,x)e^{\phi^{\mu}(x)+\psi^{\mu}(y)}f_{0}(x)f_{\mu}(y),\qquad x,y\in\mathbb{R}^{d}.$
An easy argument of separation of variables then yields that
$\rho^{\mu}(y)=e^{\psi^{\mu}(y)}f_{\mu}(y);\quad\rho_{0}(x)=e^{\phi^{\mu}(x)}f_{0}(x),\qquad x,y\in\mathbb{R}^{d}.$ (5.12)
@@
Recall the fundamental estimate (3.2) and the definition of ${\textbf{c}}(\cdot,\cdot)$, it is readily seen that $\xi(y)\sim\lambda|y|^{2}$, as $y\to\infty$, for some constant $\lambda>0$ depending only on the coefficient $b(\cdot,\cdot)$ in SDE(3.1). In light of Remark 5.5, we shall now assume, without loss of generality, that in Assumption 5.4 the control function $g$ satisfies
@@
Assume that Assumptions 2.2 and 5.4 are in force. Assume further that $\{\mu_{n}\}_{n\geq 1}\subset\mathscr{E}$ and $\mu_{n}\Longrightarrow\mu$ in Prohorov metric. Then $\|\rho^{\mu_{n}}-\rho^{\mu}\|_{\mathbb{L}^{1}}=\|\Gamma_{1}(\mu_{n})-\Gamma_{1}(\mu)\|_{\mathbb{L}^{1}}\to 0$, as $n\to\infty$.
@@
Assume $\{\mu_{n}\}_{n\geq 1}\subset\mathscr{E}$, and $\mu_{n}\Longrightarrow\mu$, in Prohorov metric. By Lemma 5.6-(ii), $\{\mu_{n}\}$ is uniformly integrable in $\mathbb{L}^{2}$, thanks to Assumption 5.4, and thus by the relationship between Wasserstein distance and Prohorov metric (see, [65, Theorem 7.12]), we have $W_{2}(\mu_{n},\mu)\to 0$, as $n\to\infty$. Thus, if follows from (5.13) that $\|\psi^{\mu_{n}}-\psi^{\mu}\|_{\infty}\to 0$, as $n\to\infty$.
@@
$\rho^{\mu_{n}}(y)=e^{\psi^{\mu_{n}}(y)}f_{\mu_{n}}(y),\quad\rho^{\mu}(y)=e^{\psi^{\mu}(y)}f_{{\mu_{\rm tar}}}(y),\qquad x,y\in\mathbb{R}^{d}.$
@@
To this end, we first recall that $\|\psi^{\mu_{n}}-\psi^{\mu}\|_{\infty}\to 0$, as $n\to\infty$. Hence there exists $N>0$, such that $\psi^{\mu_{n}}(y)\leq\psi^{\mu}(y)+1$, for all $y\in\mathbb{R}^{d}$, whenever $n\geq N$. Thus, for $n\geq N$, we have
@@
Finally, since $I_{n}^{2}(y)\leq 2\eta(y)$, and $f_{\mu_{n}}\to f_{\mu}$ in $\mathbb{L}^{1}(\mathbb{R}^{d})$, thanks to Lemma 5.6-(iii), we can apply Dominated Convergence again to get $I_{n}^{2}(\cdot)$ converges to 0 in $\mathbb{L}^{1}(\mathbb{R}^{2})$, as $n\to\infty$, proving the proposition. ∎
@@ 6 Existence of optimal control and convergence for general $\mu_{\rm ini}$}
In this section, we shall extend the results of §3 and show that the Problem 2.4 has solution for each $k\in\mathbb{N}$ when $\mu_{\rm ini}$} is an arbitrary distribution in $\mathscr{P}_{2}(\mathbb{R}^{d})$. To be more precise, for fixed $k\in\mathbb{N}$, let $J^{k}(\alpha)$ be the cost functional in Problem 2.4. Applying Lemma 5.3, for any $\mu\in\mathscr{D}$, we have
@@
Before we proceed, let us first introduce some notations. First, for any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$, we denote $f_{\mu}$ to be its density function, whenever exists. In particular, we define $f_{0}=f_{{\mu_{\rm ini}}}$. To be consistent with the previous discussions, we recall the mapping $(\nu_{0},\nu_{T})={\cal T}({\mu_{\rm ini}},\mu)$ for ${\mu_{\rm ini}},\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$, and for fixed $\mu_{\rm ini}$}, we denote $f_{\nu_{0}}=\rho_{0}$ and $f_{\nu_{T}}=\rho^{\mu}=\Gamma_{1}(\mu)$. Furthermore, from (5.1) we see that, for any ${\mu_{\rm ini}},\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$
$\frac{{\mu_{\rm ini}}(dx)}{\nu_{0}(dx)}=\frac{f_{0}(x)}{\rho_{0}(x)}=\int_{\mathbb{R}^{d}}P(T,y;0,x)\rho^{\mu}(y)dy,\quad\frac{\mu(dy)}{\nu_{T}(dy)}=\frac{f_{\mu}(y)}{\rho^{\mu}(y)}=\int_{\mathbb{R}^{d}}p(T,y;0,x)\rho_{0}(x)dx.$
In other words, we can write
$\left\{\begin{array}[]{lll}\displaystyle\rho_{0}(x)=\frac{f_{0}(x)}{\int_{\mathbb{R}^{d}}P(T,y;0,x)\rho^{\mu}(y)dy}=\frac{f_{0}(x)}{h^{\mu}(0,x)}\vskip 6.0pt plus 2.0pt minus 2.0pt,\\ \displaystyle\rho^{\mu}(y)=\frac{f_{\mu}(y)}{\int_{\mathbb{R}^{d}}p(T,y;0,x)\rho_{0}(x)dx}=\frac{f_{\mu}(y)}{\int_{\mathbb{R}^{d}}p(T,y;0,x)\frac{f_{0}(x)}{h^{\mu}(0,x)}dx}.\end{array}\right$. (6.4)
@@
Finally, we define $\Gamma=\Gamma_{2}\circ\Gamma_{1}:\mathscr{P}_{2}(\mathbb{R}^{d})\mapsto\mathscr{P}_{2}(\mathbb{R}^{d})$, and we shall argue that the mapping $\Gamma$ has a fixed point $\widehat{\mu}\in\mathscr{E}$, where $\mathscr{E}$ is defined by (5.5). Clearly, if $\Gamma(\widehat{\mu})=\widehat{\mu}$, then we can still define $\widehat{\alpha}=\alpha^{\widehat{\mu}}$, and by Lemma 5.3 we have $\mathbb{P}_{X^{\alpha^{\widehat{\mu}}}}=\widehat{\mu}$. Thus by (6.5), for any $\alpha\in\mathbb{L}^{2}_{\mathbb{F}^{0}}([0,T])$ we have
@@
To show that the minimizers $\mu^{\prime}_{n}\Rightarrow\mu^{\prime}$, we shall invoke the notion of $\Gamma$-convergence (cf. [19]). To be more precise, a sequence $\{F_{n}\}_{n\geq 1}$ is said to $\Gamma$-converge to $F$ as $n\to\infty$ if
@@
Now, note that $G$ is convex and $\mathscr{E}$ is compact under Prohorov metric, we see that both $\{F^{k}_{n}\}$ and $F^{k}$ are coercive (in the sense that there exists minimizing sequence in $\mathscr{E}\subset\mathscr{P}_{2}(\mathbb{R}^{d})$). Thus, in light of the $\Gamma$-convergence result (see [19, Theorem 7.1]), in order to show $\mu_{n}^{\prime}\Rightarrow\mu^{\prime}$, it suffices to check that $\{F^{k}_{n}\}$ $\Gamma$-converges to $F^{k}$, for each $k$. To see this, for any $\{\bar{\mu}_{n}\}\subset\mathscr{E}$, such that $\bar{\mu}_{n}\Rightarrow\bar{\mu}$, by Lemma 5.6-(iii), we have $f_{\bar{\mu}_{n}}\to f_{\bar{\mu}}$ in $\mathbb{L}^{1}$, and therefore by Dominated Convergence,

### v2

@@ 5 Stability of the Solutions to the SBP
We note that all the results in the previous section are based on an important assumption: ${\mu_{\rm ini}}=\delta_{x_{0}}$, for some $x_{0}\in\mathbb{R}^{d}$. In this and the next section, we shall extend the results to more general initial condition ${\mu_{\rm ini}}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ with density $f_{\rm ini}$, and establish a similar rate of convergence.
@@
For any ${\mu_{\rm ini}},\mu\in\mathscr{P}(\mathbb{R}^{d})$, there exists a unique pair of $\sigma$-finite measures $\nu_{0},\nu_{T}\in\mathscr{M}(\mathbb{R}^{d})$ such that the measure $\pi$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ defined by
@@
has marginals $\mu_{\rm ini}$} and $\mu$. Furthermore, $\nu_{T}$ and $\mu$ (resp. $\nu_{0}$ and $\mu_{\rm ini}$}) are mutually absolutely continuous, denoted by $\nu_{T}\simeq\mu$ (resp. $\nu_{0}\simeq{\mu_{\rm ini}}$). ∎
Fix ${\mu_{\rm ini}}\in\mathscr{P}(\mathbb{R}^{d})$ apriori. Let us denote a (well-defined) mapping ${\cal T}:\mathscr{P}_{2}(\mathbb{R}^{d})\to\mathscr{M}(\mathbb{R}^{d})\times\mathscr{M}(\mathbb{R}^{d})$ by ${\cal T}(\mu)=(\nu_{0},\nu_{T})$. Note that in Proposition 5.1 the measures $(\nu_{0},\nu_{T})$ are only $\sigma$-finite in general, to facilitate our discussion, we shall consider, the following set:
$\mathscr{D}_{{\mu_{\rm ini}}}:=\{\mu\in\mathscr{P}_{2}(\mathbb{R}^{d}):{\cal T}(\mu)\ll Leb(\cdot);\,{\cal T}(\mu)(\mathbb{R}^{d}\times\mathbb{R}^{d})<\infty\}.$
Here $Leb(\cdot)$ denotes the Lebesgue measure on $\mathbb{R}^{d}\times\mathbb{R}^{d}$.
We note that if $\mu\in\mathscr{D}_{{\mu_{\rm ini}}}$ and $(\nu_{0},\nu_{T})={\cal T}(\mu)$, then $\nu_{T}$ must have a density function, which we shall denote by $\rho^{\mu}\in\mathbb{L}^{1}(\mathbb{R}^{d})$. Moreover, we define an operator $S:\mathscr{P}_{2}(\mathbb{R}^{d})\to\mathscr{P}_{2}(\mathbb{R}^{d})$ by
@@
Moreover, for a fixed ${\mu_{\rm ini}}\in\mathscr{P}_{2}(\mathbb{R}^{d})$, we define $h^{\mu}(t,x)=\int_{\mathbb{R}^{d}}p(T,z;t,x)\rho^{\mu}(z)dz$. Then, we have the following result.
@@
Let $\mu\in{\mathscr{D}_{{\mu_{\rm ini}}}}$. Then, for any $\{\alpha_{t}\}\subset\mathbb{L}^{2}_{\mathbb{F}^{0}}([0,T];\mathbb{R}^{d})$, it holds that
@@
From Proposition 5.2 and Lemma 5.3 we see that the density function $\rho^{\mu}$ plays an important role in the structure of the solution of SBP. We shall be particularly interested in the continuous dependence of $\rho^{\mu}:=\Gamma_{1}(\mu)$ on $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$, which we shall refer to as the Stability of the SBP, borrowing the well-known concept of the SBP theory (cf. e.g., [46, 22, 10]).
@@
(i) There exists a constant $C>0$, independent of $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$, such that $|\log f_{\mu}(x)|\leq C|x|^{2}$, $x\in\mathbb{R}^{d}$;
(ii) There exists a function $g\in\mathbb{L}^{2}(\mathbb{R}^{d};(0,1])$, with $\int_{\mathbb{R}^{d}}|x|^{2}g(x)ds<\infty$, and a constant $K>0$, independent of $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$, such that $\Big\|\frac{f_{\mu}}{g^{2}}\Big\|_{\infty}\leq K$.
@@
We note that Assumption 5.4-(i) provides a uniform lower bound for the densities $f_{\mu}\in{\cal K}$, whereas Assumption 5.4-(ii) controls the decay of $f_{\mu}\in{\cal K}$ as $x\sim\infty$, hence mutually non-inclusive. A typical example of the function $g$ is $e^{-c|x|^{2}}$, $x\in\mathbb{R}^{d}$, $c>0$. In fact, in light of the estimate (3.3), such a property holds essentially for all transition probabilities of diffusion processes. ∎
@@
$f^{m}_{\mu_{n}}(x)=[\varphi^{m}*f_{\mu_{n}}](x)=\int_{\mathbb{R}^{d}}\varphi^{m}(x-z)f_{\mu_{n}}(z)dz;\quad f^{m}_{\mu}(x)=[\varphi^{m}*f_{\mu}](x),\quad x\in\mathbb{R}^{d}.$
@@
$\leq$ $K\int_{\mathbb{R}^{d}}|\varphi^{m}(z^{\prime}+y)-\varphi^{m}(z^{\prime})|dz^{\prime}\to 0,\quad\mbox{as $y\to 0$.}$
@@
We are now ready to study our main stability result. More precisely, we shall argue that the mapping $\Gamma_{1}:\mathscr{P}_{2}(\mathbb{R}^{d})\to\mathbb{L}^{1}(\mathbb{R}^{d})$ is continuous. That is, that $\mu_{n}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ weakly converges to some $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$ in Prohorov metric would imply that $\rho^{\mu_{n}}$ converges to $\rho^{\mu}$ in $\mathbb{L}^{1}(\mathbb{R}^{d})$. Such a result, to the best of our knowledge, is novel in the literature.
To simplify our discussion, in what follows, we assume $T=1$. Recall that $\pi$ in (5.1) has marginals $\mu_{\rm ini}$} and $\mu$, and in what follows, we shall assume that $\mu_{\rm ini}$} is fixed and $\mu\in\mathscr{E}$. Let us now consider the following entropic optimal transport problem:
@@
$\widehat{\pi}(dxdy)=p(1,y;0,x)e^{\phi^{\mu}(x)+\psi^{\mu}(y)}f_{\rm ini}(x)f_{\mu}(y)dxdy,$
where $\phi^{\mu}$ and $\psi^{\mu}$ are the Schrödinger potentials, and we shall enforce the symmetric normalization so that they satisfy (5.11). Since $\widehat{\pi}$ has the marginals $\mu_{\rm ini}$} and $\mu$, by the uniqueness of $(\nu_{0},\nu_{T})={\cal T}(\mu)$, in Proposition 5.1, we can conclude that
$p(1,y;0,x)\rho_{0}(x)\rho^{\mu}(y)=p(1,y;0,x)e^{\phi^{\mu}(x)+\psi^{\mu}(y)}f_{\rm ini}(x)f_{\mu}(y),\qquad x,y\in\mathbb{R}^{d},$
where $\rho_{0}$ is the density function of $\nu_{0}$. An easy argument of separation of variables then yields that
$\rho^{\mu}(y)=e^{\psi^{\mu}(y)}f_{\mu}(y);\quad\rho_{0}(x)=e^{\phi^{\mu}(x)}f_{\rm ini}(x),\qquad x,y\in\mathbb{R}^{d}.$ (5.12)
@@
Recall the fundamental estimate (3.3) and the definition of ${\textbf{c}}(\cdot,\cdot)$, it is readily seen that $|\xi(y)|\sim\lambda^{\prime}|y|^{2}$, as $y\to\infty$, for some constant $\lambda^{\prime}>0$ depending only on the coefficient $b(\cdot,\cdot)$ in SDE(3.1). In light of Remark 5.5 and [46, Lemma 2.1], in what follows we shall impose some explicit assumptions on the Schrödinger potentials $\{\psi^{\mu}(\cdot)\}_{\mu\in\mathscr{E}}$ and their upper bound $\xi(\cdot)$ in (5.14), which are in line with Assumption 5.4.
Assumption 5.8.
(i) There exists a constant $C>0$, such that for any $\mu\in\mathscr{E}$, it holds that
$|\psi^{\mu}(y)|\leq C|y|^{2},\quad y\in\mathbb{R}^{d}.$ (5.15)
(ii) In Assumption 5.4, the control function $g$ satisfies
@@
Assume that Assumptions 2.2 and 5.4 are in force. Assume further that $\{\mu_{n}\}_{n\geq 1}\subset\mathscr{E}$ and $\mu_{n}\Rightarrow\mu$ in Prohorov metric. Then $\|\rho^{\mu_{n}}-\rho^{\mu}\|_{\mathbb{L}^{1}}=\|\Gamma_{1}(\mu_{n})-\Gamma_{1}(\mu)\|_{\mathbb{L}^{1}}\to 0$, as $n\to\infty$.
@@
Assume $\{\mu_{n}\}_{n\geq 1}\subset\mathscr{E}$, and $\mu_{n}\Rightarrow\mu$, in Prohorov metric. By Lemma 5.6-(ii), $\{\mu_{n}\}$ is uniformly integrable in $\mathbb{L}^{2}$, thanks to Assumption 5.4, and thus by the relationship between Wasserstein distance and Prohorov metric (see, [65, Theorem 7.12]), we have $W_{2}(\mu_{n},\mu)\to 0$, as $n\to\infty$. Thus, if follows from (5.13) that $\|\psi^{\mu_{n}}-\psi^{\mu}\|_{\mathbb{L}^{\infty}}\to 0$, as $n\to\infty$.
@@
$\rho^{\mu_{n}}(y)=e^{\psi^{\mu_{n}}(y)}f_{\mu_{n}}(y),\quad\rho^{\mu}(y)=e^{\psi^{\mu}(y)}f_{\mu}(y),\qquad x,y\in\mathbb{R}^{d}.$
@@
To this end, we first recall that $\|\psi^{\mu_{n}}-\psi^{\mu}\|_{\mathbb{L}^{\infty}}\to 0$, as $n\to\infty$. Hence there exists $N>0$, such that $\psi^{\mu_{n}}(y)\leq\psi^{\mu}(y)+1$, for all $y\in\mathbb{R}^{d}$, whenever $n\geq N$. Thus, for $n\geq N$, we have
@@
Finally, since $I_{n}^{2}(y)\leq 2K\eta(y)$, and $f_{\mu_{n}}\to f_{\mu}$ in $\mathbb{L}^{1}(\mathbb{R}^{d})$, thanks to Lemma 5.6-(iii), we can apply Dominated Convergence again to get $I_{n}^{2}(\cdot)$ converges to 0 in $\mathbb{L}^{1}(\mathbb{R}^{2})$, as $n\to\infty$, proving the proposition. ∎
@@ 6 Existence of optimal control and convergence for general $\mu_{\rm ini}$}
In this section, we shall extend the results of Section 3 and show that the Problem 2.4 has solution for each $k\in\mathbb{N}$ when $\mu_{\rm ini}$} is an arbitrary distribution with density $f_{\rm ini}$ in $\mathscr{P}_{2}(\mathbb{R}^{d})$. To be more precise, for fixed $k\in\mathbb{N}$, let $J^{k}(\alpha)$ be the cost functional in Problem 2.4. Applying Lemma 5.3, for any $\mu\in\mathscr{D}_{\mu_{\rm ini}}$, we have
@@
Finally, we define $\Gamma=\Gamma_{2}\circ\Gamma_{1}:\mathscr{P}_{2}(\mathbb{R}^{d})\to\mathscr{P}_{2}(\mathbb{R}^{d})$, and we shall argue that the mapping $\Gamma$ has a fixed point $\widehat{\mu}\in\mathscr{E}$, where $\mathscr{E}$ is defined by (5.5). Clearly, if $\Gamma(\widehat{\mu})=\widehat{\mu}$, then we can still define $\widehat{\alpha}=\alpha^{\widehat{\mu}}$, and by Lemma 5.3 we have $\mathbb{P}_{X_{T}^{\alpha^{\widehat{\mu}}}}=\widehat{\mu}$. Thus by (6.2), for any $\alpha\in\mathbb{L}^{2}_{\mathbb{F}^{0}}([0,T];\mathbb{R}^{d})$ we have
@@
To show that the minimizers $\mu^{\prime}_{n}\Rightarrow\mu^{\prime}$, we shall invoke the notion of $\Gamma$-convergence (cf. [19]). To be more precise, to prove the sequence $\{F^{k}_{n}\}_{n\geq 1}$ $\Gamma$-converge to $F^{k}$ as $n\to\infty$, it is sufficient to show
@@
Now, note that $G$ is convex and $\mathscr{E}$ is compact under Prohorov metric, we see that both $\{F^{k}_{n}\}$ and $F^{k}$ are coercive (in the sense that there exists minimizing sequence in $\mathscr{E}\subset\mathscr{P}_{2}(\mathbb{R}^{d})$). Thus, in light of the $\Gamma$-convergence result (see [19, Theorem 7.1]), in order to show $\mu_{n}^{\prime}\Rightarrow\mu^{\prime}$, it suffices to check that $\{F^{k}_{n}\}$ $\Gamma$-converges to $F^{k}$, $k\in\mathbb{N}$. To see this, let $\{\bar{\mu}_{n}\}\subset\mathscr{E}$ such that $\bar{\mu}_{n}\Rightarrow\bar{\mu}$, we have